OpenStudy (kayley01):

Events A and B are mutually exclusive, with P(A) = 0.36 and P(B) = 0.05. What is P(A or B)? A. 0.018 B. 0.31 C. 0.41 D. 0.86

OpenStudy (anonymous):

\[P \left( A \cup B \right)=P \left( A \right)+P \left( B \right)\]

OpenStudy (kayley01):

P(0.36) + P(0.05)